### 05 Mar 2026 v3

@@ Abstract
BHs with integrable singularities (IS) have gained attention because, unlike usual regular BHs, they avoid a potentially unstable de Sitter core and the presence of an internal horizon that breaks predictability, while exhibiting finite tidal forces that allow nondestructive radial infall. First, we present a new BH solution sourced by a string fluid (FS) exhibiting an IS. We introduce an energy density profile obtained by screening the cloud of strings (CS) density within an FS framework, leading to finite conserved energy, unlike the CS case where it diverges. On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild exterior BH has recently gained attention Ovalle (2024); Maeda (2025). This is achieved by matching an interior region to the Schwarzschild exterior at the event horizon. Motivated by the variety of singular BH solutions in the literature, we establish the conditions that an interior region with an IS must satisfy to represent the interior of a generic exterior BH solution, with Schwarzschild as only a particular case. We derive the junction conditions between the interior and exterior regions, showing that they ensure temperature continuity at the interface, while discontinuities in tangential pressure signal phase transitions. We examine an interior described by CS and FS, matched to a Reissner–Nordström exterior.
@@ I Introduction
On the other hand, String theory posits that the fundamental constituents of the universe are one-dimensional objects called strings Younesizadeh and Jokar (2022). The dynamics of a single 1-brane string can be extended to an ensemble Yang et al. (2023), forming a cloud of strings. Embedding such a cloud of strings modifies the spacetime geometry, producing black hole solutions surrounded by cloud of strings Letelier (1979), and several recent studies have analyzed black holes interacting with string clouds Zafar et al. (2025); Javed et al. (2026); Gogoi et al. (2025); Ahmed et al. (2025); Cai et al. (2025). A generalization of this model includes pressure or tension in the energy–momentum tensor, giving it a structure similar to a perfect fluid, i.e., a fluid of strings Letelier (1983). In this connection, recently an equation of state $\rho(r)=\alpha(r)p(r)$ was proposed Nunes dos Santos (2025) for a fluid of strings, leading to a regular black hole solution.
@@
Following Ref. Aoki et al. (2021), where the volume integral $\int_{0}^{\infty}4\pi r^{2}T^{0}{}_{0},dr$ is interpreted as the conserved energy associated with a timelike Killing vector, we note that this integral diverges for the standard string-cloud energy density. Motivated by this, as the first problem in this work we introduce a modified version of the string-cloud energy density within the fluid-of-strings framework. This modification can be interpreted as a geometrical screening of the original distribution. As a result, we show that the volume integral converges to a finite value equal to $M$, which can be interpreted as the conserved energy charge. In this way, we provide a new fluid-of-strings black hole, constructed through the equation of state $\rho(r)=\alpha(r)p(r)$. Furthermore, we will show that the Ricci scalar near the origin behaves in a way consistent with an integrable singularity, and that the new solution is also free of a potentially unstable inner horizon.
@@
On the other hand, the Schwarzschild solution, which according to Birkhoff’s theorem Hawking and Ellis (2023) is the unique spherically symmetric vacuum solution of Einstein’s equations, possesses a strong central singularity. In this paragraph, we point out some ideas discussed in Ref. Ovalle (2024). The Schwarzschild black hole features an event horizon, which can be viewed as a membrane of radius $h$ separating two causally disconnected space-time regions: an interior and an exterior one. The interior region, defined by $r<h$, contains the central singularity shielded by the event horizon. In this domain, the metric function $f(r)$ becomes negative, causing the radial and temporal coordinates to interchange their roles and thereby revealing the dynamical character of the interior geometry. As a consequence, gravitational collapse toward the central singularity becomes unavoidable, in agreement with Penrose’s theorem Penrose (1965). Conversely, an observer in the exterior region has no causal access to the interior of the black hole: once the boundary at $r=h$ is crossed, it is impossible to return or communicate with the exterior. In this line of research, remarkably, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior has recently gained attention Ovalle (2024); Maeda (2025). This configuration is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon, without introducing a lightlike thin shell. Reference Ovalle (2024) examines whether an alternative structure, beyond a pointlike mass, can give rise to the Schwarzschild black hole region for $r\geq h$. In particular, the authors aim to describe an interior configuration associated with this region while keeping the exterior Schwarzschild geometry unchanged and ensuring that tidal forces remain finite everywhere. In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core endowed with a potentially unstable inner horizon. Reference Maeda (2025) constructs a so-called fake Schwarzschild solution, observationally indistinguishable from the Schwarzschild black hole, by matching a dynamical spherically symmetric interior containing matter to the Schwarzschild exterior at the event horizon without introducing a lightlike thin shell. It is shown that this configuration can be obtained from a perfect-fluid solution satisfying a linear equation of state of the type considered in Ref. Maeda (2023), analytically continued inside the horizon.
It is well known that a wide class of black hole solutions studied in the literature are supported by matter sources that inevitably lead to a central singularity, just to mention a few examples: black holes with cosmic-void density profiles Lustosa et al. (2025), black holes surrounded by dark matter with dark energy profiles Xu et al. (2018), hairy black holes Ovalle et al. (2021), and quintessential black holes Kiselev (2003); Estrada et al. (2025). Furthermore, electrically charged BH, exemplified by the Reissner–Nordström solution, also arise from electromagnetic sources in the energy–momentum tensor and are characterized by the presence of an inner horizon. Motivated by the above considerations, and in analogy with Ref. Ovalle (2024), we address as a second problem the existence of an internal geometry inside the event horizon of a generic black hole spacetime, keeping the external geometry completely general, with the Schwarzschild solution appearing only as a particular case. We focus on scenarios in which the internal structure is endowed with an integrable singularity, in order to avoid the emergence of a potentially unstable de Sitter core and the presence of an inner horizon, since predictability breaks down after crossing the inner horizon. We investigate the possible nature of matter sources capable of supporting such an internal geometry with an integrable singularity, specifically those associated with clouds of strings and with the new string fluid introduced in this work. To do this, we define the conditions that the internal geometry must satisfy in order to represent an integrable singularity, as well as the junction conditions between generic internal and external black hole geometries. Furthermore, we examine the connection between the Israel–Darmois Israel (1966) junction conditions and black hole thermodynamic properties, and explore their relation to the existence of phase transitions. As a test case, we study the Reissner–Nordström solution as an internal geometry, considering clouds of strings and the new string fluid as the first and second examples, respectively.
@@
II A new Fluid of strings black hole solution with an integrable singularity
@@
In this section, we first present a new string fluid solution, showing that it can represent a black hole solution over the entire domain $r\in[0,\infty)$.Below, in line with this work, we characterize it as the interior region of a Reissner–Nordström black hole. Reference Letelier (1983) proposes a generalization of the cloud of strings model by incorporating pressure or tension into the energy–momentum tensor. This model is the so-called fluid of strings (fs). See Appendix B for further details. Due to the symmetries of the spacetime and in connection with the structure of the energy–momentum tensor (70), its general form is given by $T^{t}_{t}=T^{r}_{r}=-\rho^{(fs)}\,,\,T^{\theta}_{\theta}=T^{\phi}_{\phi}=p^{(fs)}$. It is of interest to establish an equation of state for the string fluid. In Ref. Nunes dos Santos (2025), an equation of state was proposed in which the energy density and the pressure are related through the form $\rho(r)^{(fs)}=\alpha(r)\cdot p(r)^{(fs)}$. In this reference, an equation of state with $\alpha(r)$ was proposed such that the solution of Einstein’s equations represents a regular string-fluid geometry at the origin. Within this framework, the energy–momentum tensor takes the form
@@
As mentioned above, in this section the radial coordinate extends over the entire domain $r\in[0,\infty)$ and, consequently, there is an asymptotically flat boundary. As indicated in Ref. Nunes dos Santos (2025), within this framework the Einstein equations lead to a solution of the form:
@@ II.2 Absence of an inner horizon
First, we note that function (9) behaves near the origin as $f_{(fs)}|_{r\sim 0}\sim 1-2M/b$, with $2M/b>1$ in order for the metric signature to be +,,,+,-,-,- near the origin. Moreover, $|1-\frac{2M}{b}|$ is sufficiently large and dominant in the vicinity of the origin, so as to avoid the appearance of a zero of the function $f(r)_{(fs)}$ near the origin. We observe the numerical behavior and the absence of a potentially unstable inner horizon in Fig. 1 . Consequently, our solution lacks the presence of an unstable de Sitter core, unlike other black hole solutions sourced by matter.
@@
In this way, in this section we have presented a new string fluid solution that represents a geometry with an integrable singularity and the absence of a potentially unstable inner horizon. As mentioned above, in this section we analyze a black hole solution over the entire domain $r\in[0,\infty)$. Below, in line with this work, we characterize it as the interior region of a Reissner–Nordström black hole.
III Black Hole Solutions with an Interior Geometry Featuring an Integrable Singularity
@@ III.1 Description of the interior spacetime :
where we observe that $\rho=-p_{r}$. The conservation equation reads:
@@ III.2 Description of the exterior spacetime
The geometry is such that $f_{E}(h)=0$. Since this region represents the exterior of the black hole, the signature of the metric tensor (25) is ,+,+,+-,+,+,+, and therefore $f_{E}(r)\geq 0$ in this region. We consider the case without a cosmological constant, with an energy–momentum tensor of the form
@@ IV INTERIOR REGION AS AN INTEGRABLE SINGULARITY
In this section, we state the properties that a generic interior solution must satisfy in order to represent an integrable singularity near the origin. In the following sections, we analyze the specific cases in which the interior regions correspond, respectively, to a string cloud and to the new string-fluid solution provided in Section II.
@@
    where $a>1$ and |1a||1-a| is sufficiently large and dominant compared to the term $\mathcal{O}(r^{n})$ near the origin, in order to avoid a zero of the function $f(r)$ in the vicinity of the origin. As mentioned above, the absence of zeros in this interval implies the absence of a potentially unstable inner horizon. Although the Ricci scalar has a singularity $\sim r^{-2}$ at the origin, the equations of motion (28) are integrable, being free of singularities, and lead to a finite solution $f(r)$ at the origin. This situation can be referred to as an integrable singularity. Under these assumptions, one can also observe that the Kretschmann scalar exhibits a singularity at $r=0$.
@@
    As mentioned, $f(r)$ is an increasing function from its minimum value $f(r=0)=f_{\min}<0$ up to $f(h)=0$. In this way, $df/dr>0$ for $r\in[0,h]$, with the temperature $T=\left.df/dr\right|_{h^{-}}=\left.df_{E}/dr\right|_{h^{+}}$.
@@ V Junction conditions
In Section  III, we described our geometric setup to represent both the interior and exterior regions of a generic black hole geometry. In Section IV, we formulated the conditions that an interior region must satisfy in order to represent an integrable singularity near the origin, while in Section V we presented the junction conditions between the interior and exterior regions. However, the criteria introduced above must be tested in explicit black hole geometries. In this spirit, we analyze two illustrative examples. In the present section, we study a Reissner–Nordström exterior with a cloud of strings as the interior region. In the following section, we consider a Reissner–Nordström exterior with the new string-fluid solution provided in Section  II.
@@
Thus, we note that it must be satisfied that $M>Q\Rightarrow M^{2}-Q^{2}>0$. As we have mentioned, in our setup, the exterior geometry extends over $r\in[h,r_{b}]$, with the boundary corresponding to the radial coordinate approaching infinity, since the solution is asymptotically flat. Therefore, we are interested in the aforementioned range of the radial coordinate. The temperature is given by:
@@ VI.1 Representation of a Cloud of Strings as an Interior Region with an Integrable Singularity
Thus, from equations (20),(21), (22), we note that the interior energy density and effective pressures take the following form
@@
In order to test condition IV 2 for the interior region to be represented as an integrable singularity, we note that $f(r)$ is an increasing function, starting from its minimum value f(r=0)=1a<0f(r=0)=1-a<0, with $a>1$, and reaching $f(h)=0$. Indeed, $df/dr=2r/l^{2}>0$ for $r\in[0,h]$, and therefore the function is monotonically increasing and negative throughout this interval. Therefore, we point out that the parameter $a$ must satisfy $a>1$ in order for this condition to be fulfilled.

### 09 Mar 2026 v4

@@ Abstract
Black Holes (BHs) with integrable singularities (IS) have attracted attention because, unlike usual regular BHs, they avoid a potentially unstable de Sitter core and the presence of an internal horizon that breaks predictability, while exhibiting finite tidal forces that allow nondestructive radial infall. On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild exterior BH region has recently gained attention Ovalle (2024); Maeda (2025). In this work, motivated by the existence of an inner horizon and a non integrable singularity in the Reissner Nordström solution (RN), we address the possibility that the RN BH possesses an interior region with an IS and no inner horizon, instead of a pointlike mass. We propose that the nature of this interior region is given by a cloud of strings (CS) with IS and by a new fluid of strings (FS) with IS, determining the physical properties that both must satisfy in order to support such a configuration. To this end, we establish the matching conditions between these interior regions with IS and the RN exterior BH solution. This leads to the requirement that the continuity of temperature must be satisfied at the interface, while discontinuities in the tangential pressure signal phase transitions. Thus, for both interior regions, we determine the conditions relating them to the RN exterior region that must be satisfied in order to physically represent the interior region with an IS while fulfilling the matching conditions. Furthermore, we identify the conditions under which phase transitions occur or do not occur.
@@ I Introduction
On the other hand, String theory posits that the fundamental constituents of the universe are one-dimensional objects called strings Younesizadeh and Jokar (2022). The dynamics of a single 1-brane string can be extended to an ensemble Yang et al. (2023), forming a cloud of strings (CS). Embedding such a cloud of strings modifies the spacetime geometry, producing black hole solutions surrounded by cloud of strings Letelier (1979), and several recent studies have analyzed black holes interacting with string clouds Zafar et al. (2025); Javed et al. (2026); Gogoi et al. (2025); Ahmed et al. (2025); Cai et al. (2025). A generalization of this model includes pressure or tension in the energy–momentum tensor, giving it a structure similar to a perfect fluid, i.e., a fluid of strings (FS) Letelier (1983). In this connection, recently an equation of state $\rho(r)=\alpha(r)p(r)$ was proposed Nunes dos Santos (2025) for a fluid of strings, leading to a regular black hole model.
@@
Following the definition of the conserved energy associated with a timelike Killing vector given in Ref. Aoki et al. (2021), we note that this conserved energy diverges for the standard cloud of strings energy density. Motivated by this, we first introduce in this work a modified version of the string-cloud energy density within the fluid-of-strings framework. This modification can be interpreted as a geometrical screening of the original distribution. As a result, this leads to the convergence of the conserved energy charge. In this way, we first provide a new fluid-of-strings black hole model, constructed through the equation of state $\rho(r)=\alpha(r)p(r)$. Furthermore, we show that the Ricci scalar near the origin behaves in a way consistent with an integrable singularity, and that the new FS model is also free of a potentially unstable inner horizon.
@@
On the other hand, the Schwarzschild solution, which according to Birkhoff’s theorem Hawking and Ellis (2023) is the unique spherically symmetric vacuum solution of Einstein’s equations, possesses a strong central singularity. As pointed out in reference Ovalle (2024), the Schwarzschild black hole features an event horizon, which can be viewed as a membrane of radius $h$ separating two causally disconnected space-time regions: an interior and an exterior one. The interior region, defined by $r<h$, contains the central singularity shielded by the event horizon. In this domain, the metric function $f(r)$ becomes negative, causing the radial and temporal coordinates to interchange their roles and thereby revealing the dynamical character of the interior geometry. As a consequence, gravitational collapse toward the central singularity becomes unavoidable, in agreement with Penrose’s theorem Penrose (1965). Conversely, an observer in the exterior region has no causal access to the interior of the black hole: once the boundary at $r=h$ is crossed, it is impossible to return or communicate with the exterior. In this line of research, remarkably, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior has recently gained attention Ovalle (2024); Maeda (2025). This configuration is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon, without introducing a lightlike thin shell. Reference Ovalle (2024) examines whether an alternative structure, beyond a pointlike mass, can give rise to the Schwarzschild black hole region for $r\geq h$. In particular, the authors aim to describe an interior configuration associated with this region while keeping the exterior Schwarzschild geometry unchanged and ensuring that tidal forces remain finite everywhere. In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core. Reference Maeda (2025) constructs a so-called fake Schwarzschild model, observationally indistinguishable from the Schwarzschild black hole, by matching a dynamical spherically symmetric interior containing matter to the Schwarzschild exterior at the event horizon without introducing a lightlike thin shell. It is shown that this configuration can be obtained from a perfect-fluid solution satisfying a linear equation of state of the type considered in Ref. Maeda (2023), analytically continued inside the horizon.
Reissner Nordström solution represents an electrically charged BH, arising from electromagnetic sources in the energy momentum tensor and characterized by the presence of an inner horizon and a central singularity. Motivated by the considerations described in this introduction, and by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), in this work, we study the physical and thermodynamic properties of a scenario in which the RN BH possesses an interior region with an integrable singularity and no inner horizon, instead of a pointlike mass. We investigate the possibility that the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity corresponds to those associated with clouds of strings and with the new string fluid introduced in this work.
To carry out this work, we will first define the conditions that a general internal geometry must satisfy in order to represent an integrable singularity, as well as the Israel Darmois Israel (1966) junction conditions between a general internal region and a general external BH geometry. We will then examine the connection between the junction conditions and the thermodynamic properties of BHs, and explore their relation to the existence of phase transitions. In this context, we will show that satisfying the junction conditions leads to the requirement that the continuity of temperature must hold at the interface, while discontinuities in the tangential pressure signal phase transitions. It is worth mentioning that these general conditions could also be of physical interest for future studies of other BH exterior regions, instead of RN, that possess a central singularity, for example BHs with cosmic void density profiles Lustosa et al. (2025), BHs surrounded by dark matter Xu et al. (2018), hairy BHs Ovalle et al. (2021), and quintessential BHs Kiselev (2003); Estrada et al. (2025). In this way, applying the general conditions described above, in this work we focus on the physical and thermodynamic properties of the mentioned configuration where the exterior geometry is given by the Reissner Nordström BH, while the interior geometry is described by a cloud and by a new fluid of strings with integrable singularities. Specifically, we will determine the physical properties that both the cloud of strings (CS) and the new fluid of strings (FS) must satisfy in order to represent the interior of an exterior RN BH region. Using the general analysis described above, we will determine the relations between the parameter spaces of both interior regions and the RN exterior BH for which the matching conditions are satisfied and, as a consequence, the temperature remains continuous at the interface. In addition, for both interior regions we will determine the conditions relating them to the RN exterior region that must be satisfied for phase transitions to occur or not occur.
@@
II A new Fluid of strings black hole model with an integrable singularity
@@
In this section, we first present a new string fluid model, showing that it can represent a black hole model over the entire domain $r\in[0,\infty)$. Below, in line with this work, we characterize it as the interior region of a Reissner–Nordström black hole. Reference Letelier (1983) proposes a generalization of the cloud of strings model by incorporating pressure or tension into the energy–momentum tensor. This model is the so-called fluid of strings. See Appendix B for further details. Due to the symmetries of the spacetime and in connection with the structure of the energy–momentum tensor (70), its general form is given by $T^{t}_{t}=T^{r}_{r}=-\rho^{(fs)}\,,\,T^{\theta}_{\theta}=T^{\phi}_{\phi}=p^{(fs)}$. It is of interest to establish an equation of state for the string fluid. In Ref. Nunes dos Santos (2025), an equation of state was proposed in which the energy density and the pressure are related through the form $\rho(r)^{(fs)}=\alpha(r)\cdot p(r)^{(fs)}$. In this reference, an equation of state with $\alpha(r)$ was proposed such that the model of Einstein’s equations represents a regular string-fluid geometry at the origin. Within this framework, the energy–momentum tensor takes the form
@@
As mentioned above, in this section the radial coordinate extends over the entire domain $r\in[0,\infty)$ and, consequently, there is an asymptotically flat boundary. As indicated in Ref. Nunes dos Santos (2025), within this framework the Einstein equations lead to a geometry of the form:
@@ II.2 Absence of an inner horizon
First, we note that function (9) behaves near the origin as $f_{(fs)}|_{r\sim 0}\sim 1-2M/b$, with $2M/b>1$ in order for the metric signature to be +,,,+,-,-,- near the origin. Moreover, $|1-\frac{2M}{b}|$ is sufficiently large and dominant in the vicinity of the origin, so as to avoid the appearance of a zero of the function $f(r)_{(fs)}$ near the origin. We observe the numerical behavior and the absence of a potentially unstable inner horizon in Fig. 1 . Consequently, our model lacks the presence of an unstable de Sitter core, unlike other black hole solutions sourced by matter.
@@
In this way, in this section we have presented a new string fluid model that represents a geometry with an integrable singularity and the absence of a potentially unstable inner horizon. As mentioned above, in this section we analyze a black hole model over the entire domain $r\in[0,\infty)$. Below, in line with this work, we characterize it as the interior region of a Reissner–Nordström black hole.
III Black Hole models with an Interior Geometry Featuring an Integrable Singularity
@@ III.1 Description of the interior spacetime :
where we observe that $\rho=-p_{r}$.
@@ III.2 Description of the exterior spacetime
The geometry is such that $f_{E}(h)=0$. Since this region represents the exterior of the black hole, the signature of the metric tensor (25) is ,+,+,+-,+,+,+, and therefore $f_{E}(r)\geq 0$ in this region. We consider the case without a cosmological constant in the external region, with an energy–momentum tensor of the form
@@ IV INTERIOR REGION AS AN INTEGRABLE SINGULARITY
In this section, we state the properties that a generic interior geometry must satisfy in order to represent an integrable singularity near the origin. In the following sections, we analyze the specific cases in which the interior regions correspond, respectively, to a string cloud and to the new string-fluid model provided in Section II.
@@
    where $a>1$ and |1a||1-a| is sufficiently large and dominant compared to the term $\mathcal{O}(r^{n})$ near the origin, in order to avoid a zero of the function $f(r)$ in the vicinity of the origin. As mentioned above, the absence of zeros in this interval implies the absence of a potentially unstable inner horizon. Although the Ricci scalar has a singularity $\sim r^{-2}$ at the origin, the equations of motion (28) are integrable, being free of singularities, and lead to a finite geometry $f(r)$ at the origin. This situation can be referred to as an integrable singularity. Under these assumptions, one can also observe that the Kretschmann scalar exhibits a singularity at $r=0$.
@@
    As mentioned, $f(r)$ is an increasing function from its minimum value $f(r=0)=f_{\min}<0$ up to $f(h)=0$. In this way, $df/dr>0$ for $r\in[0,h]$, with the temperature $T=\left.df/dr\right|_{h^{-}}=\left.df_{E}/dr\right|_{h^{+}}$. It is worth mentioning that this condition leads to the absence of an inner horizon.
@@ V Junction conditions
In Section  III, we described our geometric setup to represent both the interior and exterior regions of a generic black hole geometry. In Section IV, we formulated the conditions that an interior region must satisfy in order to represent an integrable singularity near the origin, while in Section V we presented the junction conditions between the interior and exterior regions. However, the criteria introduced above must be tested in explicit black hole geometries. In this spirit, motivated by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), we analyze two illustrative examples. In the present section, we study a Reissner–Nordström exterior with a cloud of strings as the interior region. In the following section, we consider a Reissner–Nordström exterior with the new string-fluid model provided in Section  II.
@@
Thus, we note that it must be satisfied that $M>Q$ for $M,Q>0$, $\Rightarrow M^{2}-Q^{2}>0$. As we have mentioned, in our setup, the exterior geometry extends over $r\in[h,r_{b}]$, with the boundary corresponding to the radial coordinate approaching infinity, since the solution is asymptotically flat. Therefore, we are interested in the aforementioned range of the radial coordinate. The temperature is given by:
@@ VI.1 Representation of a Cloud of Strings as an Interior Region with an Integrable Singularity
Thus, from equations (20),(21), (22), we note that the interior effective energy density and effective pressures take the following form
@@
In order to test conditions IV and 2 for the interior region to be represented as an integrable singularity without an inner horizon, we note that $f(r)$ is an increasing function, starting from its minimum value f(r=0)=1a<0f(r=0)=1-a<0, with $a>1$, and reaching $f(h)=0$. Indeed, $df/dr=2r/l^{2}>0$ for $r\in[0,h]$, and therefore the function is monotonically increasing and negative throughout this interval. Therefore, we point out that the parameter $a$ must satisfy $a>1$ in order for this condition to be fulfilled.